We wish to estimate what percent of adult residents in a certain county are parents. Out of 500 adult residents sampled, 320 had kids. Based on this, construct a 95% confidence interval for the proportion p of adult residents who are parents in this county.

Express your answer in tri-inequality form. Give your answers as decimals, to three places.

___< p <____ Express the same answer using the point estimate and margin of error. Give your answers as decimals, to three places.

p = ___ ± ___
Answer #1

we have given: n= 500 x=320 320 500 95% confidence interrad for Proportion (p) P=[ť 1 12612 * 84 (hp)] :: 2412 = 20.05 = [1 1= (0.64 + (1.96 * 0.0214)] = 10.64 + 0.0420) = [ 0.5970, 0.682] => 1 0.597/< p < 0.682) we can say 95 . confident proportion

> Where did you get 0.36?

josephin samking Sun, Nov 21, 2021 5:00 PM
Answer #2

where did you get 0.36?

answered by: josephin samking